OpenStudy (anonymous):

The radius of a cylinder is 3.5 ft. The height is 14 ft. Find the surface area and volume of the cylinder to the nearest tenth of a foot. Show your work

Parth (parthkohli):

\[A = 2\pi r(r + h)\]\[V = \pi r^2 h\]

Parth (parthkohli):

Work with those formulas. T_T

OpenStudy (anonymous):

ok :)

Parth (parthkohli):

do they want ya to use \(\large \frac{22}{7}\) or \(3.14\) for \(\pi\)?

OpenStudy (anonymous):

3.14

Parth (parthkohli):

OK then.